# Comparison results for Markov tree distributions

### v1

@@ Appendix B Proofs of Section 1
In this section, we provide the proof of Theorem 1.3 which is based on combining Lemma 1.2 with an extension of the conditioning argument in [58, Theorem 3.2] to a non-stationary setting and to Markovian tree structures. For working with trees, this requires additional tools, which we introduce in the sequel. It is worth noting that Proposition 1.1 serves as an extension of [58, Theorem 3.2] to a non-stationary context. This proposition emerges as a direct consequence of Theorem 1.3.
@@
In order to navigate through a tree, we define the commonly used notions of a parent, child, descendant, and ancestor relationship as follows.
@@
    The set of children of $i\in N$ is defined by $c_{i}\coloneqq\{j\in N|i=p_{j}\}\subset N.$
@@
    The set of descendants of $i\in N$ is defined by $d_{i}\coloneqq\{j\in N|\text{there is a directed path from $i$ to $j$}\}\subset N.$
@@
    The set of ancestors of $i\in N\backslash\{0\}$ is defined by $a_{i}\coloneqq\{j\in N|i\in d_{j}\}\subset N.$
  5. (v)
    The degree of a node $i\in N\backslash\{0\}$ is defined by $\operatorname{deg}(i)\coloneqq|c_{i}|+1$ and $\operatorname{deg}(0)\coloneqq|c_{0}|$.
According to its definition, a node $i\in N$ is not included in the set of its children $c_{i}$, descendants $d_{i}$, or ancestors $a_{i}$. Note that the root $0\in N$ is the only element having no parent node.
The proof of Theorem 1.3 is based on an induction, which requires a specific enumeration of the nodes of the tree. To this end, we consider the level-order traversal where the nodes of the tree are visited level by level, starting from the root. This is also known as a breadth-first search. To be precise, we say that $\pi=(\pi_{n})_{n\in N}$ is an enumeration of $N$ if it is a one-to-one map from $N$ to $N$. Further, $\pi$ is a level-order traversal if it is an enumeration such that for any pair of nodes $i$ and $j$ with $i<j$ (with respect to the order on $\mathbb{N}$), the node $\pi_{i}$ is in a lower level of the tree compared to $\pi_{j}$, i.e., for all $i,j,\in N\,,$
$i<j\quad\text{implies}\quad|p[0,\pi_{i}]|\leq|p[0,\pi_{j}]|.$ (45)
Note that this enumeration is not uniquely determined. We examplify the level order traversal in Figure 9 within a tree that comprises 8 nodes distributed across 3 levels.
@@
To prove Theorem 1.3, we rely on several lemmas presented in the sequel. The subsequent lemma shows that if any random variable is SI in its ancestor random variables in a Markov tree distributed sequence, then any random vector corresponding to descendant nodes of this node is SI in the random variable associated to this current node. In the context of Markov processes, this means that if at each time step, the present random variable is SI in the preceding variable, then all future variables are SI in the present variable, see [45, Theorem 3.2]. We define, in comparison to Definition 2.1, the undirected paths
@@
Moreover, for a subset of nodes $J\subseteq N$, we write $x_{J}\coloneqq(x_{j})_{j\in J}$ and $X_{J}\coloneqq(X_{j})_{j\in J}$ when referencing to the vector and random vector associated with these components.
@@
  • Proof:
    We prove the result by induction over $|J|$. Let $|J|=1$, $J=\{j_{1}\}\subseteq d_{i}$ for $i\in N$. Then, by assumption, $X_{k}\uparrow_{st}X_{p_{k}}$ for all $k\in p(i,j_{1}]$, see (46). We conclude by [45, Theorem 3.2] that $X_{j_{1}}\uparrow_{st}X_{i}$. For the induction step, assume that the statement is valid for all $J\subseteq d_{i}$ such that $|J|=n-1$. Suppose that $J$ has the form $J=\{j_{1},\dots,j_{n}\}\subseteq d_{i}$. The following part of the proof is divided in two cases. First, we assume that there are two nodes in $J$ which are not separated by $i$, i.e.,
@@
    Then there is a node $j_{\ell}\in J$ such that $J\cap p(i,j_{\ell})\neq\emptyset$. In particular, the node $j_{\ell}$ can be choosen such that $j_{\ell}\notin p(i,j)$ for all $j\in J$. Let $j_{k}\in J$ be the last node in $J$ in the directed path from $i$ to $j_{\ell}$, i.e.,
@@
    $=\int\underbrace{\int\phi(x_{j_{\ell}},x_{K})P^{X_{j_{\ell}}|X_{j_{m}}=x_{j_{m}}}(dx_{j_{\ell}})}_{\coloneqq\psi(x_{j_{m}},x_{K\backslash\{j_{m}\})}}P^{X_{K}|X_{i}=x_{i}}(dx_{K}),$ (48)
    where we use the Markov tree dependence for the second equality. Applying [45, Theorem 3.2] once more, we can conclude that $\psi$ is an increasing and bounded function in $x_{j_{m}}$ for all $x_{K\backslash\{j_{m}\}}\in\mathbb{R}^{n-2}$. Moreover, since $\phi$ is increasing, also the function $\psi$ is increasing in $x_{K\backslash\{j_{m}\}}$ for all $x_{j_{m}}\in\mathbb{R}$. Consequently, $\psi$ is increasing in each argument. By the induction hypothesis, we have that $X_{K}\uparrow_{st}X_{i}$ since $|K|=n-1$ and $K\subseteq d_{i}$. Hence, by the definition of the stochastic order, the function in (48) is increasing in $x_{i}$, proving that $X_{J}\uparrow_{st}X_{i}$.
In the second case we assume that $J\cap p(i,j)=\emptyset$ for all $j\in J=\{j_{1},\dots,j_{n}\}$. For $K\coloneqq J\backslash\{j_{n}\}$ we have for any $\phi\colon\mathbb{R}^{n}\to\mathbb{R}$ increasing and bounded that
@@
$\begin{split}&\int\phi(x_{J})P^{X_{J}|X_{i}=x_{i}}(dx_{J})\\ &=\int\int\phi(x_{J})P^{X_{K}|X_{i}=x_{i}}(dx_{K})P^{X_{j_{n}}|X_{i}=x_{i}}(dx_{j_{n}}).\end{split}$ (50)
@@
The following lemma, which we were unable to locate in the literature, shows that if all components $X_{i}$, $i\in N\backslash\{0\}$ of a Markov tree distributed random vector in a random vector $X$ are SI in its parents, then $X_{i}$ is SI in all of its ancestors.
@@
Note that property (51) is in general not a positive dependence concept unless $X$ follows a Markov tree distribution. For example, let $T=(N,E)$, $N=\{(1,2),(1,3)\}$. For some independet and uniformly on $[0,1]$ distributed random variables $U$ and $V$ we define $X_{1}=U$, $X_{2}=V$, $X_{3}=1-V$. Then $X=(X_{1},X_{2},X_{3})$ meets (51), but is not positive lower (upper) orthant dependent, see Definition 2.8.
However, in the context of chains, property (51) is equivalent to the CIS property, and thus is a positive dependence concept in the sense of (14).
The following lemma extends [58, Theorem 3.1] from Markov processes to Markov tree distributions. It provides sufficient condition such that an integral of a supermodular function, under a conditional distribution, yields a supermodular function dependent on both the conditioning variable and the variables that remain unintegrated. Denote by $p[i,j]:=p(i,j)\cup\{i\}\cup\{j\}$ the undirected path from $i$ to $j$ including $i$ and $j\,.$
@@
    Let $X$ be a random vector that follows a Markov tree distribution with respect to a tree $T=(N,E)$. Let $i\in N$ and $J\subseteq d_{i}$ be a finite set. Assume that $X_{j^{\prime}}\uparrow_{st}X_{j}$ for all $(j,j^{\prime})\in E$. Moreover let $K$ be an arbitrary finite index set and $\phi$ be a supermodular function on $\mathbb{R}^{1+|K|+|J|}$. Then the function $\varphi\colon\mathbb{R}^{|K|+1}\to\mathbb{R}$ given by
@@
    $(z_{k_{1}},z_{k_{2}})\mapsto\varphi(z_{i},z_{k_{1}},z_{k_{2}},z_{K\backslash\{k_{1},k_{2}\}})$ (54)
    is supermodular for all $(z_{i},z_{K\backslash\{k_{1},k_{2}\}})\in\mathbb{R}^{1+|K|-2}$. In the following we show for $k\in K$ that the map
    $(z_{i},z_{k})\mapsto\varphi(z_{i},z_{k},z_{K\backslash\{k\}})$ (55)
    is supermodular for all $z_{K\backslash\{k\}}\in\mathbb{R}^{|K|-1}$. We fix $z_{k}<z_{k}^{\prime}$. By the supermodularity of $\phi$ it follows for all $j\in J$ that the map
@@
    is increasing for all $z_{J\backslash\{j\}}\in\mathbb{R}^{|J|-1}$ and $z_{K\backslash\{k\}}\in\mathbb{R}^{|K|-1}$. Similarly, the map
@@
    is increasing for all $z_{J}\in\mathbb{R}^{|J|}$ and $z_{K\backslash\{k\}}\in\mathbb{R}^{|K|-1}$. By Lemma B.3 the assumption $X_{j^{\prime}}\uparrow_{st}X_{j}$ for all $(j,j^{\prime})\in E$ implies that $X_{j}\uparrow_{st}(X_{j^{\prime}},j^{\prime}\in a_{j})$ for all $j\in N\backslash\{0\}$. Thus, we can conclude from Lemma B.2 that $(X_{j},j\in J)\uparrow_{st}X_{i}$. Now, for $z_{i}<z_{i}^{\prime}$, we obtain that
    $\begin{split}&\varphi(z_{i},z_{k}^{\prime},z_{K\backslash\{k\}})-\varphi(z_{i},z_{k},z_{K\backslash\{k\}})\\ &\leq\int\phi(z_{i},z^{\prime}_{k},z_{K\setminus\{k\}},z_{J})P^{X_{J}|X_{i}=z^{\prime}_{i}}(dz_{J})-\int\phi(z_{i},z_{k},z_{K\backslash\{k\}},z_{J})P^{X_{J}|X_{i}=z^{\prime}_{i}}(dz_{J})\\ &\leq\varphi(z^{\prime}_{i},z_{k}^{\prime},z_{K\backslash\{k\}})-\varphi(z^{\prime}_{i},z_{k},z_{K\backslash\{k\}}),\end{split}$ (58)
    where the first inequality follows by $(X_{j},j\in J)\uparrow_{st}X_{i}$ and the increasingness of (56) for $j\in J\,.$ The second inequality holds true because the function in (57) is increasing. Now, (LABEL:eq:_lemma4_monotonicity) implies that
    $\varphi(z^{\prime}_{i},z_{k}^{\prime},z_{K\backslash\{k\}})+\varphi(z_{i},z_{k},z_{K\backslash\{k\}})\geq\varphi(z^{\prime}_{i},z_{k},z_{K\backslash\{k\}})+\varphi(z_{i},z^{\prime}_{k},z_{K\backslash\{k\}})$
@@
The following proposition is a version of Theorem 1.3 under additional SI assumptions. This result is used for the proof of Theorem 1.3 where we show that several SI conditions along a specified path can be omitted. We briefly outline the idea behind the proof of this proposition.
We initiate with the distribution of $X=(X_{n})_{n\in N}$ and replace the dependence specifications of $X$ for the specifications of $Y$ iteratively over all star structures within the tree $T=(N,E)$ using a conditioning argument and applying Lemma 1.2 on the supermodular comparison of star structures. More precisely, for the induction, we navigate through the tree using a level-order traversal enumeration as described in (45). A disintegration argument allows us to reduce the proof of the statement in each step to the setting of Lemma 1.2. Finally, by the transitivity of the supermodular order, we obtain a supermodular comparison between $X$ and $Y$.
@@
Let $X=(X_{n})_{n\in N}$ and $Y=(Y_{n})_{n\in N}$ be sequences of random variables that follow a Markov tree distribution with respect to a tree $T=(N,E)$. Assume that
@@
  • Proof:
    By the definitions of the stochastic orderings for stochastic processes (see Definition 2.7), we can assume that $d+1:=|N|<\infty$ and $N=\{0,\ldots,d\}$. We only show the case that $X\leq_{sm}Y$ if $(X_{i},X_{j})\leq_{sm}(Y_{i},Y_{j})$ for all $(i,j)\in E$. In the case where $(X_{i},X_{j})\geq_{sm}(Y_{i},Y_{j})$ for all $(i,j)\in E$, it follows analogously that $X\geq_{sm}Y$ by reversing all inequality signs.
    Denote by $F=(F_{0},\dots,F_{d})$ the vector of univariate marginal distribution functions of $X$ and $Y$. Note that $F_{X_{n}}=F_{Y_{n}}$ for all $n\in N$ because $(X_{i},X_{j})\leq_{sm}(Y_{i},Y_{j})$ implies $X_{i}\stackrel{{\scriptstyle\mathrm{d}}}{{=}}Y_{i}$ and $X_{j}\stackrel{{\scriptstyle\mathrm{d}}}{{=}}Y_{j}\,.$ Moreover, for $(i,j)\in E$, denote by $B_{(i,j)}$ and $C_{(i,j)}$ a copula for $(X_{i},X_{j})$ and $(Y_{i},Y_{j})$, respectively, and define $B=(B_{e})_{e\in E}$ and $C=(C_{e})_{e\in E}$. It follows that $X\sim\mathcal{M}(F,B,T)$ and $Y\sim\mathcal{M}(F,C,T)$, see Proposition 2.5. Let $(\pi_{n})_{n=1,\dots,d}$ be a level-order traversal of $T$ as defined in (45). For $i\in N$, define the set of edges connecting the node $i$ with its children by $[i,c_{i}]\coloneqq\{(i,j)|j\in c_{i}\}\subseteq E$. Then $[i,c_{i}]$ has a star-like structure. Note that if $i\in L$, then $c_{i}=\emptyset$ and thus $[i,c_{i}]=\emptyset$. For $k=0,\dots,d-1$, let $D^{k}=(D^{k}_{e})_{e\in E}$ be the set of bivariate copulas defined by
    $D_{e}^{k}\coloneqq\begin{cases}C_{e}&\text{if }e\in[\pi_{n},c_{\pi_{n}}]\text{ for }1\leq n\leq k,\\ B_{e}&\text{if }e\in[\pi_{n},c_{\pi_{n}}]\text{ for }k<n\leq d.\\ \end{cases}$ (59)
@@
[…]
    $E[\phi(Z^{(k-1)})]\leq E[\phi(Z^{(k)})]\,,$ (61)
    see [87, Theorem 3.9.14]. Recall that $d_{i}$ denotes the descendants of node $i\,,$ see Definition B.1. Then we define the function $\varphi^{(0)}$ on $\mathbb{R}^{|d_{i}|+1}$ by
@@
    $\varphi^{(\ell)}(z_{i},z_{c_{i}},z_{d_{j_{\ell+1}}},\dots,z_{d_{j_{m}}})$
    $\coloneqq\int\varphi^{(\ell-1)}(z_{i},z_{c_{i}},z_{d_{j_{\ell}}},\dots,z_{d_{j_{m}}})P^{Z_{d_{j_{\ell}}}^{(k)}|Z^{(k)}_{j_{\ell}}=z_{j_{\ell}}}(dz_{d_{j_{\ell}}}),$
@@
To prove supermodularity of $\varphi^{(0)}$, we will apply Lemma B.4. Recall that $p[0,i]$ denotes the path from 00 up to and including node $i$. From the assumptions (ii) and (i) we know, in particular, that
@@
where, for a set $M\,,$ the Cartesian product $M\times M$ is denoted by $M^{2}\,.$ Consider the subtree $T_{0}=(N_{0},E_{0})$ with nodes $N_{0}:=N\backslash d_{i}$ and edges $E_{0}:=E\cap N_{0}^{2}$. The nodes $N_{0}$ can be relabeled such that $i$ becomes the new root in $T_{0}$. By (62) the Markov tree distributed sequence $(Z^{(k)}_{n})_{n\in N_{0}}$ (with respect to the tree $T_{0}$) satisfies the assumptions of Lemma B.4 (ii) and by setting $K=d_{i}$ and $J=N_{0}\backslash\{i\}$ (here $d_{i}$ still denotes the set of descendents of $i$ with respect to $T$) we obtain the supermodularity of $\varphi^{(0)}$.
@@
Hence, we obtain supermodularity of $\varphi^{(\ell)}$ by Lemma B.4 (ii) by setting $K=\{i\}\cup(c_{i}\backslash\{j_{\ell}\})\cup d_{j_{\ell+1}}\cup\dots\cup d_{j_{m}}$ and $J=d_{j_{\ell}}$. and thus obtain the supermodularity of $\varphi^{(\ell)}$ by Lemma B.4 (ii). In particular, we obtain that $\varphi^{(m)}$ is supermodular.
@@
and
$\varphi^{(\ell)}(z_{i},z_{c_{i}},z_{d_{j_{l+1}}},\dots,z_{d_{j_{m}}})$
$=\int\varphi^{(\ell-1)}(z_{i},z_{c_{i}},z_{d_{j_{\ell}}},\dots,z_{d_{j_{m}}})P^{Z_{d_{j_{\ell}}}^{(k)}|Z^{(k)}_{i}=z_{i},Z^{(k)}_{c_{i}}=z_{c_{i}},Z^{(k)}_{d_{j_{\ell+1}}}=z_{d_{j_{\ell+1}}},\dots,Z^{(k)}_{d_{j_{m}}}=z_{d_{j_{m}}}}(dz_{d_{j_{\ell}}})$ (67)
$=\int\varphi^{(\ell-1)}(z_{i},z_{c_{i}},z_{d_{j_{\ell}}},\dots,z_{d_{j_{m}}})P^{Z_{d_{j_{\ell}}}^{(k-1)}|Z^{(k-1)}_{i}=z_{i},Z^{(k-1)}_{c_{i}}=z_{c_{i}},Z^{(k-1)}_{d_{j_{\ell+1}}}=z_{d_{y_{\ell+1}}},\dots,Z^{(k-1)}_{d_{j_{m}}}=z_{d_{j_{m}}}}(dz_{d_{j_{\ell}}})$

### v3

@@ B.1 Proofs of Section 1
In this section, we provide the proof of Theorem 1.3 which combines Lemma 1.2 with an extension of the conditioning argument in [38, Theorem 3.2] to a non-stationary setting and to Markovian tree structures. Working with trees requires additional tools, which we introduce in the sequel.
@@
To navigate through a tree, we define the commonly used terms for parent, child, descendant, and ancestor relationships as follows.
@@
    The set of children of $i\in N$ is defined by $c_{i}\coloneqq\{j\in N|i=p_{j}\}.$
@@
    The set of descendants of $i\in N$ is defined by $d_{i}\coloneqq\{j\in N|\text{ there exists}$ $\text{a directed path from $i$ to $j$}\}.$
@@
    The set of ancestors of $i\in N\backslash\{0\}$ is defined by $a_{i}\coloneqq\{j\in N|i\in d_{j}\}.$
Recall that the degree of a node $i\in N\backslash\{0\}$ is given by $\operatorname{deg}(i)=|c_{i}|+1$ and $\operatorname{deg}(0)=|c_{0}|$. According to its definition, a node $i\in N$ is not included in the set of its children $c_{i}$, descendants $d_{i}$, or ancestors $a_{i}$. Note that the root $0\in N$ is the only node without parent.
The proof of Theorem 1.3 is based on induction, which requires a specific enumeration of the nodes of the tree. To this end, we consider the level-order traversal, where the nodes of the tree are visited level by level, starting from the root. This is also known as a breadth-first search. Note that in the proofs, we only consider trees with finitely many children, as the ordering results need to be proved only for the finite-dimensional marginal distributions. Formally, we say that $\pi=(\pi_{n})_{n\in N}$ is an enumeration of $N$ if it is a one-to-one map from $N$ to $N$. Further, $\pi$ is a level-order traversal if it is an enumeration such that for any pair of nodes $i$ and $j$ with $i<j$ (with respect to the order on $\mathbb{N}$), the node $\pi_{i}$ is in a lower or the same level of the tree compared to $\pi_{j}$, i.e., for all $i,j,\in N,$
$i<j\quad\text{implies}\quad|p[0,\pi_{i}]|\leq|p[0,\pi_{j}]|,$ (42)
where $p[0,\pi_{i}]$ denotes an undirected path from the root to $\pi$; see (43) below. Note that this enumeration is not uniquely determined. We illustrate the level-order traversal in Figure 9 using a tree with $8$ nodes distributed across $3$ levels.
@@
To prove Theorem 1.3, we rely on several lemmas presented in the sequel. The following lemma for Markov tree distributions states that, if any random variable is SI in its ancestor random variables, then any random vector corresponding to descendant nodes is SI in the random variable associated with the current node. In the context of Markov processes, this means that if, at each time step, the present random variable is SI in the preceding variable, then all future variables are SI in the present variable; see [33, Theorem 3.2]. We define, similarly to the undirected path $p(i,j)$ introduced in Section 2.1.1, the undirected paths
@@
Moreover, for a subset of nodes $J\subseteq N$, we write $x_{J}\coloneqq(x_{j})_{j\in J}$ and $X_{J}\coloneqq(X_{j})_{j\in J}$ to refer to the vector and random vector associated with these components.
@@
We prove the result by induction over $|J|$. Let $|J|=1$, $J=\{j_{1}\}\subseteq d_{i}$ for $i\in N$. Then, by assumption, $X_{k}\uparrow_{st}X_{p_{k}}$ for all $k\in p(i,j_{1}]$; see (43). We conclude by [33, Theorem 3.2] that $X_{j_{1}}\uparrow_{st}X_{i}$. For the induction step, assume that the statement holds for all $J\subseteq d_{i}$ such that |J|=n1|J|=n-1. Suppose that $J$ has the form $J=\{j_{1},\dots,j_{n}\}\subseteq d_{i}$. The following part of the proof is divided in two cases. First, we assume that there are two nodes in $J$ which are not separated by $i$, i.e.,
@@
Then there exists a node $j_{\ell}\in J$ such that $J\cap p(i,j_{\ell})\neq\emptyset$. In particular, the node $j_{\ell}$ can be chosen such that $j_{\ell}\notin p(i,j)$ for all $j\in J$. Let $j_{k}\in J$ be the last node in $J$ in the directed path from $i$ to $j_{\ell}$, i.e.,
@@
[…]
In the second case we assume that $J\cap p(i,j)=\emptyset$ for all $j\in J=\{j_{1},\dots,j_{n}\}$. For $K\coloneqq J\backslash\{j_{n}\}$ we have for any increasing and bounded $\phi\colon\mathbb{R}^{n}\to\mathbb{R}$ that
@@
$\begin{split}&\int\phi(x_{J})P^{X_{J}|X_{i}=x_{i}}(dx_{J})=\int\int\phi(x_{J})P^{X_{K}|X_{i}=x_{i}}(dx_{K})P^{X_{j_{n}}|X_{i}=x_{i}}(dx_{j_{n}}).\end{split}$ (47)
@@
The following simple lemma, which we were unable to locate in the literature, shows that if all components $X_{i}$, $i\in N\backslash\{0\}$ of a Markov tree distributed random vector are SI in its parents, then $X_{i}$ is SI in all of its ancestors.
@@
The following lemma extends [38, Theorem 3.1] from Markov processes to Markov tree distributions. It provides sufficient conditions under which an integral of a supermodular function, with respect to a conditional distribution, results in a supermodular function that depends on both the conditioning variable and the variables that remain unintegrated. Recall that $p[i,j]$ denotes the undirected path from $i$ to $j$, including $i$ and $j.$
@@
    Let $X$ be a random vector that follows a Markov tree distribution with respect to a tree $T=(N,E)$. Let $i\in N$ and let $J\subseteq d_{i}$ be a finite set. Assume that $X_{j^{\prime}}\uparrow_{st}X_{j}$ for all $(j,j^{\prime})\in E$. Moreover let $K$ be an arbitrary finite index set and $\phi$ be a supermodular function on $\mathbb{R}^{1+|K|+|J|}$. Then the function $\varphi\colon\mathbb{R}^{|K|+1}\to\mathbb{R}$ given by
@@
[…]
@@
[…]
@@
[…]
where the first inequality follows by $(X_{j},j\in J)\uparrow_{st}X_{i}$ and by the increasingness of (53) for $j\in J.$ The second inequality holds true because the function in (54) is increasing. Now, (LABEL:eq:_lemma4_monotonicity) implies
[…]
@@
The following proposition is a version of Theorem 1.3 under additional SI assumptions. We briefly outline the idea behind the proof of this proposition. We start with the distribution of $X=(X_{n})_{n\in N}$ and iteratively replace the dependence specifications of $X$ with those of $Y$ over all star structures within the tree $T=(N,E)$. This procedure uses a conditioning argument and applies Lemma 1.2 on the supermodular comparison of star structures. More precisely, for the induction, we navigate through the tree using a level-order traversal enumeration as described in (42). A disintegration argument reduces the proof at each step to the setting of Lemma 1.2. Finally, by leveraging the transitivity of the supermodular order, we obtain a supermodular comparison between $X$ and $Y$. In the proof of Theorem 1.3, we further show that several SI conditions along a specified path can be omitted.
@@
Let $X=(X_{n})_{n\in N}$ and $Y=(Y_{n})_{n\in N}$ be sequences of random variables each following a Markov tree distribution with respect to a tree $T=(N,E)$. Let
@@
By definition of the stochastic orderings for stochastic processes (see Definition 2.5), we can assume that $d+1:=|N|<\infty$ and $N=\{0,\ldots,d\}$. We only show the case that $X\leq_{sm}Y$ if $(X_{i},X_{j})\leq_{sm}(Y_{i},Y_{j})$ for all $(i,j)\in E$. In the case where $(X_{i},X_{j})\geq_{sm}(Y_{i},Y_{j})$ for all $(i,j)\in E$, it follows analogously that $X\geq_{sm}Y$ by reversing all inequality signs.
Denote by $F=(F_{0},\dots,F_{d})$ the vector of univariate marginal distribution functions of $X$ and $Y$. Note that $F_{X_{n}}=F_{Y_{n}}$ for all $n\in N$ because $(X_{i},X_{j})\leq_{sm}(Y_{i},Y_{j})$ implies $X_{i}\stackrel{{\scriptstyle\mathrm{d}}}{{=}}Y_{i}$ and $X_{j}\stackrel{{\scriptstyle\mathrm{d}}}{{=}}Y_{j}.$ Moreover, for $(i,j)\in E$, denote by $B_{(i,j)}$ and $C_{(i,j)}$ a copula for $(X_{i},X_{j})$ and $(Y_{i},Y_{j})$, respectively, and define $B=(B_{e})_{e\in E}$ and $C=(C_{e})_{e\in E}$. It follows that $X\sim\mathcal{M}(F,B,T)$ and $Y\sim\mathcal{M}(F,C,T)$; see Proposition 2.4. Let $(\pi_{n})_{n=1,\dots,d}$ be a level-order traversal of $T$ as defined in (42). For $i\in N$, define the set of edges connecting the node $i$ with its children by $i,c_{i}]\coloneqq\{(i,j)|j\in c_{i}\}\subseteq E$. Then $i,c_{i}]$ has a star-like structure. Note that if $i\in L$, then $c_{i}=\emptyset$ and thus $i,c_{i}]=\emptyset$. For $k=0,\dots,d-1$, let $D^{k}=(D^{k}_{e})_{e\in E}$ be the set of bivariate copulas defined by
@@
i.e., the specifications of the edges in the first $k$ stars in level-order traversal are replaced by the bivariate copula specifications of $Y$, and the remaining edges are specified by the bivariate copulas associated to $X.$ For $k\in\{0,\ldots,d-1\},$ consider the random vector $Z^{(k)}\sim\mathcal{M}(F,T,D^{k})$. Note that $D^{0}=B$ and $D^{d-1}=C$, which implies that $Z^{(0)}\overset{\text{d}}{=}X$ and $Z^{(d)}\overset{\text{d}}{=}Y$. We show that
@@
Then the statement follows from the transitivity of the supermodular order.To prove (57), fix $k\in\{1,\dots,d-1\}$. If $\pi_{k}\in L$ we have that $D^{k-1}=D^{k}$ and thus $Z^{(k-1)}\overset{\text{d}}{=}Z^{(k)}$ which trivially implies $Z^{(k-1)}\leq_{sm}Z^{(k)}$. If $\pi_{k}\notin L$, define $i\coloneqq\pi_{k}\in N$ and denote by $\{j_{1},j_{2},\dots,j_{m}\}\coloneqq c_{i}$ an enumeration of the children $c_{i}$ of node $i.$ Note that, by the definition of the bivariate copula sequences in (56), $D^{k-1}$ and $D^{k}$ can only differ in the copulas which correspond to the edges $i,c_{i}]$ between node $i$ and its children, i.e., $D_{e}^{k-1}=D_{e}^{k}$ for all $e\in E\backslash[i,c_{i}]$. For proving (57), it is sufficient to show for any bounded supermodular function $\phi\colon\mathbb{R}^{d+1}\to\mathbb{R}$ that
$E[\phi(Z^{(k-1)})]\leq E[\phi(Z^{(k)})]\,;$ (58)
see [61, Theorem 3.9.14]. Recall that $d_{i}$ denotes the descendants of node i;i\,; see Definition B.1. Then we define the function $\varphi^{(0)}$ on $\mathbb{R}^{|d_{i}|+1}$ by
@@
$\varphi^{(\ell)}(z_{i},z_{c_{i}},z_{d_{j_{\ell+1}}},\dots,z_{d_{j_{m}}})\coloneqq\int\varphi^{(\ell-1)}(z_{i},z_{c_{i}},z_{d_{j_{\ell}}},\dots,z_{d_{j_{m}}})P^{Z_{d_{j_{\ell}}}^{(k)}|Z^{(k)}_{j_{\ell}}=z_{j_{\ell}}}(dz_{d_{j_{\ell}}}),$
@@
To prove supermodularity of $\varphi^{(0)}$, we will apply Lemma B.4. Recall that $p[0,i]$ denotes the path from 0 up to and including node $i$. From the assumptions (ii) and (i) we know, in particular, that
@@
To prove the supermodularity of $\varphi^{(\ell)}$, $\ell\in\{1,\ldots,m\},$ we assume by induction that $\varphi^{(\ell-1)}$ is supermodular. By definition of $Z^{(k)}$ and assumption (i) we have that
@@
Hence, we obtain supermodularity of $\varphi^{(\ell)}$ by Lemma B.4 (ii) by setting $K=\{i\}\cup(c_{i}\backslash\{j_{\ell}\})\cup d_{j_{\ell+1}}\cup\dots\cup d_{j_{m}}$ and $J=d_{j_{\ell}}$. In particular, we obtain that $\varphi^{(m)}$ is supermodular.
@@
and, for $\mathbb{z}=(z_{i},z_{c_{i}},z_{d_{j_{\ell}}},\dots,z_{d_{j_{m}}}),$
$\quad\quad\varphi^{(\ell)}(z_{i},z_{c_{i}},z_{d_{j_{l+1}}},\dots,z_{d_{j_{m}}})$
$=\int\varphi^{(\ell-1)}(\mathbb{z})P^{Z_{d_{j_{\ell}}}^{(k)}|Z^{(k)}_{i}=z_{i},Z^{(k)}_{c_{i}}=z_{c_{i}},Z^{(k)}_{d_{j_{\ell+1}}}=z_{d_{j_{\ell+1}}},\dots,Z^{(k)}_{d_{j_{m}}}=z_{d_{j_{m}}}}(dz_{d_{j_{\ell}}})$ (64)
$=\int\varphi^{(\ell-1)}(\mathbb{z})P^{Z_{d_{j_{\ell}}}^{(k-1)}|Z^{(k-1)}_{i}=z_{i},Z^{(k-1)}_{c_{i}}=z_{c_{i}},Z^{(k-1)}_{d_{j_{\ell+1}}}=z_{d_{y_{\ell+1}}},\dots,Z^{(k-1)}_{d_{j_{m}}}=z_{d_{j_{m}}}}(dz_{d_{j_{\ell}}})$